OpenStudy (anonymous):

The velocity of sound in air is given by the equation where v is the velocity in meters per second and t is the temperature in degrees Celsius. Find the temperature when the velocity of sound in air is 369 meters per second. Round to the nearest degree. A. 507ºC B. 6,535ºC C. 7,081ºC D. 67ºC

OpenStudy (anonymous):

the equation is v=2o square root of 273 t

OpenStudy (anonymous):

20 sorry

OpenStudy (unklerhaukus):

\[v=20\sqrt{273}T\]

OpenStudy (anonymous):

idk how to solve that equation

OpenStudy (unklerhaukus):

did i type it correctly?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

you are asked to solve for T when v=369 \[v=20\sqrt{273}T\] divide both sides of the equation by 20 sqrt{273} \[\frac{v}{20\sqrt{273}}=T\] now replace v with 369

OpenStudy (unklerhaukus):

something is not right, is the t meant to be inside the square root?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so how do i get the degrees in celcius

OpenStudy (unklerhaukus):

CAN you draw the equation i still think something is missing,, like a plus or minus

OpenStudy (anonymous):

sorry i forgot you add the t to the square root

OpenStudy (unklerhaukus):

\[v=20 + \sqrt{273 t}\]

OpenStudy (anonymous):

no\[v=20\sqrt{273}+t\]

OpenStudy (anonymous):

does that help

OpenStudy (unklerhaukus):

\[v=20\sqrt{273+t}\]\[\frac{v}{20}=\sqrt{273+t}\]\[\frac{v^2}{400}=273+t\]\[\frac{v^2}{400}-273=t\]\[\frac{369^2}{400}-273=t\]

OpenStudy (unklerhaukus):

This works